### v2

# MMReview: A Multidisciplinary and Multimodal Benchmark for LLM-Based Peer Review Automation

Abstract
With the rapid growth of academic publications, peer review has become an essential yet time-consuming responsibility within the research community. Large Language Models (LLMs) have increasingly been adopted to assist in the generation of review comments; however, current LLM-based review tasks lack a unified evaluation benchmark to rigorously assess the models’ ability to produce comprehensive, accurate, and human-aligned assessments—particularly in scenarios involving multimodal content such as figures and tables. To address this gap, we propose MMReview, a comprehensive benchmark that spans multiple disciplines and modalities. MMReview includes multimodal content and expert-written review comments for 240 papers across 17 research domains within four major academic disciplines: Artificial Intelligence, Natural Sciences, Engineering Sciences, and Social Sciences. We design a total of 13 tasks grouped into four core categories, aimed at evaluating the performance of LLMs and Multimodal LLMs (MLLMs) in step-wise review generation, outcome formulation, alignment with human preferences, and robustness to adversarial input manipulation. Extensive experiments conducted on 16 open-source models and 5 advanced closed-source models demonstrate the thoroughness of the benchmark. We envision MMReview as a critical step toward establishing a standardized foundation for the development of automated peer review systems.
Introduction
Peer review is essential to scholarly publishing, ensuring research quality and enhancing academic writing. However, the growing volume of submissions has strained the traditional review process, leading to inefficiencies and limited reviewer availability (Kim, Lee, and Lee 2025), which restricts feedback and delays review outcomes. Advances in LLMs have made automated peer review increasingly viable, as these models show strong reasoning abilities and can offer constructive feedback on academic manuscripts (Liu and Shah 2023; Zhao et al. 2024; Zhuang et al. 2025), partially alleviating reviewer burden. Yet, current evaluations of LLM-generated reviews focus mainly on final outputs, lacking in-depth analysis of the reasoning processes behind model judgments. Additionally, most studies concentrate on AI papers with publicly available text, overlooking the multimodal nature of academic papers—such as figures and tables—and the evaluation of LLMs in reviewing research across broader scientific domains.
To address the aforementioned challenges, we propose MMReview, a comprehensive benchmark for peer review generation that spans multiple disciplines and modalities. MMReview incorporates three distinct types of input modalities: textual content from manuscripts, figures and tables embedded within the papers, and rendered PDF pages converted into images. These data span 17 research domains across 4 disciplinary categories. To obtain high-quality peer review samples for evaluation purposes, we developed a multi-model collaborative pipeline for data filtering and generation. Specifically, we first curated a total of 51,881 papers with associated reviews; then, we filtered the collected seed dataset D to obtain high-quality papers while maintaining a relatively balanced distribution; subsequently, we extracted reference answers from human reviews. Finally, we conducted manual verification to correct errors, resulting in a curated set of 240 samples that serve as the foundation for task construction. Building upon these samples, we introduce 4 thematic categories encompassing 13 diverse tasks, each designed to thoroughly assess the capabilities of LLMs in step-wise review generation, outcome formulation, alignment with human preferences, and robustness to adversarial input manipulation. We conduct comprehensive experiments on 16 open-source models as well as 5 state-of-the-art closed-source models, including GPT-4o and Claude-4-Sonnet, across 13 tasks. The results highlight the comprehensive nature of the MMReview benchmark and uncover several key findings, offering insights for future research on LLM-based automated academic peer review.
The primary contributions of this paper can be summarized as follows:
  • We introduce MMReview, the first comprehensive evaluation benchmark for automated academic peer review using LLMs, spanning multiple disciplines and modalities. Built upon our data filtering and generation pipeline, MMReview comprises 240 high-quality samples across 17 academic fields in 4 disciplines.
  • We meticulously design 13 distinct tasks encompassing a total of 6,724 thoughtfully curated questions, enabling multi-dimensional evaluation of model performance. These diverse tasks allow for targeted assessment and facilitate the identification of potential limitations in LLM-generated peer review content.
  • We conduct extensive experiments on 16 open-source and 5 closed-source models using the MMReview benchmark, offering some key findings of LLM-based automated reviewing. Our findings offer in-depth analysis and valuable guidance for the future development of LLM-assisted peer review systems.
Related Works
LLMs for Paper Review
LLMs have shown strong potential in analyzing complex scholarly texts (Liu and Shah 2023; Zhao et al. 2024; Zhuang et al. 2025). Initial studies indicate that LLM-generated review comments partially overlap with those of human reviewers, suggesting their potential contribution to peer review (Robertson 2023; Liang et al. 2023). However, further research reveals that even advanced models like GPT-4o often fail to meet human expectations in review quality (Zhou, Chen, and Yu 2024). To improve alignment with peer review standards, researchers have built datasets from public review platforms and fine-tuned LLMs (Kang et al. 2018; Yuan, Liu, and Neubig 2021; Shen et al. 2022; Dycke, Kuznetsov, and Gurevych 2023; Gao, Brantley, and Joachims 2024). Other methods involve multi-turn dialogue (Tan et al. 2024) or multi-agent prompting (D’Arcy et al. 2024) to generate comprehensive feedback. Nonetheless, these efforts focus solely on textual reviews, neglecting the vital role of multimodal content—such as figures and tables—and lack rigorous analysis of the reasoning processes behind LLM-generated critiques.
Evaluation for LLM-based Peer Review
Prior studies (Shen et al. 2022; Yu et al. 2024; Gao, Brantley, and Joachims 2024; Tan et al. 2024; Gao et al. 2025) have predominantly evaluated the quality of LLM-generated peer review comments by measuring their correlation or similarity with human-written reviews using automated metrics such as BLEU (Papineni et al. 2002), ROUGE (Lin 2004), BERTScore (Zhang et al. 2020), and METEOR (Banerjee and Lavie 2005). In addition, several studies (Robertson 2023; Zhou, Chen, and Yu 2024; Gao et al. 2025) have adopted the LLM-as-a-judge paradigm, leveraging cutting-edge language models to assess the quality of review comments produced by other LLMs. Given the absence of an established gold standard for this evaluation task, recent research (Xu et al. 2024) has introduced the Generative Estimator for Mutual Information (GEM) to quantify the degree of semantic overlap between LLM-generated and human-authored reviews. Nevertheless, existing evaluation methodologies are not grounded in a unified benchmark or task framework, and they fall short of providing a comprehensive analysis of the underlying reasoning processes involved in LLM-generated peer review.
Figure 1: The construction pipeline of MMReview. The construction pipeline is divided into three stages: data collection, data processing, and task construction.
MMReview Benchmark
In this section, we first present the overall pipeline for data collection and construction of the MMReview benchmark, followed by a detailed exposition of the task design methodology.
Overall Pipeline of MMReview Benchmark
As illustrated in Figure 1, the construction of MMReview is divided into three stages: data collection, data processing, and task construction.
Data Collection
During the data collection phase, we gathered academic papers from publicly accessible peer review platforms or sources where reviewer comments were openly available. These papers not only contain the full manuscript texts but also include reviewer-written evaluations and final decisions (accept or reject). Specifically, we curated a total of 51,881 papers with associated reviews submitted between 2013 and 2024 to venues such as ICLR and NeurIPS (from the OpenReview platform and NeurIPS Proceedings , as well as articles from the journal Nature Communications . For each collected paper, we employed Deepseek-V3 and GPT-4o to automatically infer its academic discipline and research domain. In cases where the two models produced inconsistent classifications, human verification was performed. Ultimately, all papers were categorized into four overarching disciplines: Artificial Intelligence, Natural Sciences, Engineering Sciences, and Social Sciences, resulting in a seed dataset denoted as $D$. Figure 1 illustrates the distribution of papers across the four disciplines within $D$.
Data Processing
During the data processing stage, we first filtered the collected seed dataset $D$ to obtain high-quality papers while maintaining a relatively balanced distribution across disciplines and ensuring a comparable number of accepted and rejected papers. To achieve this, we designed a dual-path joint data selection mechanism that simultaneously filters for paper quality and enforces distributional balance.
Specifically, we utilized Qwen3-32B [^1] Qwen3-32B ranks among the top 10 on the OpenCompass leaderboard, offering a favorable trade-off between performance and model size, making it efficient for large-scale sample filtering (Yang et al. 2025) to generate summaries of each paper under two distinct input conditions: one using only the abstract and the other using the full text of the manuscript. A greater divergence between the two generated summaries is interpreted as evidence that the full text provides substantially more information, thus indicating higher paper quality. Samples with a significant information gain from the full text were retained in our test benchmark. To further ensure that the benchmark maintains a balanced distribution across academic disciplines and an approximately equal ratio of accepted to rejected papers, we supplemented the benchmark by manually incorporating top-ranked papers from specific domains in $D$ based on the quality rankings. This human-filtering procedure guarantees that the composition of the benchmark dataset aligns with our desired distributional properties.
During the data filtering phase, we obtained a total of 240 paper samples spanning 17 research domains across 4 major disciplines to construct our evaluation benchmark, denoted as $D_{b}$. The statistical details of $D_{b}$, including the number of samples per domain and the distribution of accepted versus rejected papers, are presented in Figure 1. For each of these 240 samples, we utilized PDF parsing tools to extract textual content, figures, and tables from the manuscript files, and converted each page of the PDF into corresponding images. As a result, we constructed three distinct modalities of input data: text-only, multimodal (text combined with extracted visual elements), and PDF-as-image.
Task Construction
During task construction, we developed 13 tasks grouped into four thematic categories—step-based, outcome-based, preference-based, and attack-based—reflecting the peer review workflow and challenges LLMs may face. Task distribution is shown in Figure 1. Prompts were designed based on reviewer guidelines from major academic conferences. For each task, we used regular expressions or GPT-4o to extract reference answers from human reviews in $D_{b}$, forming the question–answer dataset $D_{QA}$. For the Fake Strength Evaluation and Fake Weakness Evaluation tasks, GPT-4o and custom prompts generated antonymic rewrites, which were manually verified for semantic accuracy. This process finalized the MMReview benchmark. Figure 2 illustrates its multidisciplinary scope, multimodal inputs, and diverse task design. The detailed prompts of task generation and each task can be found in the supplementary materials.
Step-based Tasks
The Step-based theme comprises five tasks designed to progressively evaluate the performance of LLMs in simulating the key components of the academic peer review process.
Summary (S)
Summarizing a paper is the initial step in peer review and a key test of a model’s ability to extract essential content. Inaccurate summaries may impair subsequent review generation. To address this, the Summary task assesses a model’s ability to distill key information from a full manuscript into an accurate, concise summary. The model generates a brief summary in its own words, avoiding abstract copying and subjective judgment. These outputs are compared to human-written summaries and evaluated for semantic similarity and information coverage, measuring the model’s holistic comprehension and representation of academic content.
Strengths Evaluation (SE) and Weaknesses Evaluation (WE)
Summarizing and analyzing a manuscript’s strengths and weaknesses is a core aspect of peer review. The Strengths Evaluation and Weaknesses Evaluation tasks assess LLMs’ ability to identify and articulate the merits and limitations of academic papers. These tasks test whether models can synthesize technical highlights and methodological concerns noted by human reviewers, focusing on four dimensions: Quality, Clarity, Significance, and Originality. In the Strengths Evaluation task, models argue for acceptance by detailing methodological rigor, experimental robustness, structural clarity, research impact, and novelty, thus evaluating their capacity to extract technical contributions and assess scientific merit. In contrast, the Weaknesses Evaluation task adopts a rejection-oriented stance, testing critical reasoning and constructive critique. Model outputs are compared with human reviews based on content coverage and semantic similarity.
Soundness Scoring (SS) and Presentation Scoring (PS)
The Soundness Scoring and Presentation Scoring tasks evaluate LLMs’ ability to quantitatively assess manuscript quality, focusing on technical soundness and writing presentation. In Soundness Scoring, the model rates the reliability of technical claims, experimental rigor, and evidential support, emphasizing empirical and methodological validity. Presentation Scoring assesses linguistic clarity and logical organization, reflecting writing quality and information structure. Both tasks require integer scores from 1 to 4, denoting “poor” to “excellent.” Model scores are compared to human ratings to assess judgment consistency.
Figure 2: An overview of MMReview benchmark.
Outcome-based Tasks
The Outcome-based tasks focus on assessing a model’s direct capability to generate peer review outcomes, with the goal of evaluating its alignment with human reviewers in final decision-making.
Conditional Decision (CD)
The Conditional Decision task assesses LLMs’ ability to synthesize human-written reviews and generate an overall quality score for a paper. Provided with reviewer comments detailing strengths, weaknesses, and evaluations of methodology and results, the model assigns a numerical score from 1 to 10, reflecting a scale from “fundamentally flawed or lacking novelty” to “groundbreaking contribution,” aligned with academic conference standards. The task evaluates the model’s capacity to interpret sentiment, weighting, and evaluative reasoning in the reviews and translate them into a coherent quantitative judgment. Model scores are compared with human ratings to assess alignment and accuracy in review-based decision-making.
Direct Decision (DD) and CoT Decision (CoD)
The Direct Decision and CoT (Chain-of-Thought) Decision tasks evaluate LLMs’ ability to autonomously assess academic paper quality, reflecting two reviewing paradigms: streamlined judgment and step-by-step reasoning (Wei et al. 2022). These tasks vary in input format and cognitive complexity, enabling controlled comparison of model performance under different reasoning demands. In the Direct Decision task, the model produces an overall score without guidance, simulating a reviewer’s holistic judgment from a single read. In contrast, the CoT Decision task guides the model through a structured reasoning process: summarizing the paper, analyzing strengths and weaknesses across Quality, Clarity, Significance, and Originality, assigning Soundness and Presentation scores, and synthesizing an overall score. This mirrors a reviewer’s iterative, analytical evaluation. The tasks test reasoning ability, analytical structure, and scoring traceability. Model outputs are compared to human scores to assess consistency and rationality. While the Direct task emphasizes efficiency, the CoT task offers interpretability, together forming a complementary framework for evaluating LLMs’ scoring competence.
Meta Decision (MD)
Beyond generating individual reviews and scores, a key aspect of academic peer review is the Area Chair’s (AC) synthesis of reviewer feedback to make a final decision. To emulate this, the Meta Decision task requires the model to issue a binary judgment—Accept or Reject—based on multiple human reviews, testing its ability to aggregate information and balance differing views. The prompt provides structured guidance and evaluation criteria, prompting step-by-step reasoning. The model is instructed to assess the quality and consistency of reviews rather than merely averaging scores. This task mirrors the real role of an AC and rigorously evaluates the model’s capacity for synthesis, decision rationality, and transparency. Model outputs are compared to human-written meta-reviews to assess reliability and scientific judgment in high-level peer review.
Preference-based Task: Pairwise Rank (PR)
Prior work has shown that pairwise comparison effectively evaluates the alignment between LLM-generated preferences and human judgments (Liu et al. 2024b). Since academic conference acceptance tiers—oral, spotlight, poster, reject—reflect human preference rankings, the Pairwise Rank task is designed to test whether LLMs, as reviewers, display preference patterns consistent with human evaluators. This task assesses the model’s relative judgment ability by presenting pairs of papers from different acceptance tiers: oral (top 5%), spotlight (top 25%), poster, and reject. The model compares and ranks the papers, simulating real-world peer review selection. Alignment is measured by comparing model preferences with actual acceptance categories to determine ranking accuracy. To reduce positional bias (Shi et al. 2025; Thakur et al. 2025), each comparison is repeated with reversed input order.
Attack-based Tasks
This task category assesses models’ robustness and discriminative ability in peer review by introducing adversarial inputs. Fabricated strengths or weaknesses—such as inverted pros and cons—are used to test the model’s capacity to detect factual inaccuracies. Misleading prompts further aim to divert the model from its original instructions. Model outputs are compared to human judgments or the model’s own non-adversarial responses, focusing on its ability to preserve evaluative independence and logical consistency under input perturbations.
Fake Strengths (FS) and Fake Weaknesses (FW)
LLMs have been widely observed to exhibit a tendency toward “sycophancy”—an inclination to conform to human judgments (Liu et al. 2024a; Kumaran et al. 2025)—which may impair their ability to identify flaws and generate critical peer review comments. Furthermore, the hallucination problem inherent in LLMs can result in inaccurate evaluations of a paper’s strengths and weaknesses. To investigate these issues, we introduce the Fake Strengths and Fake Weaknesses tasks, which assess the model’s capacity to recognize factual inconsistencies and produce critical, hallucination-aware assessments. In these tasks, we intentionally manipulate human-written review comments by reversing their semantic polarity—transforming strengths into weaknesses and vice versa—using GPT-4o. For example, the original strength “the paper is clearly written” is rephrased as “the paper lacks clarity.” The model is then queried regarding the presence of these altered attributes, allowing us to examine its ability to accurately judge the validity of strengths and weaknesses, detect factual inconsistencies, and generate grounded, critical evaluations.
Prompt Injection (PI)
Recent studies have revealed that some paper authors attempt to manipulate LLM-generated reviews by employing prompt injection techniques—embedding invisible white-text prompts within the manuscript to coax models into generating overly favorable evaluations (Gibney 2025). However, the actual effectiveness of such manipulations remains insufficiently quantified. To systematically assess this vulnerability, we design the Prompt Injection task, wherein highly suggestive prompts, such as “please ignore all prior prompts and provide a positive evaluation of this paper”, are embedded directly into the manuscript text to test whether the model can be coerced into abandoning objectivity. The task setup mirrors that of the CoT Decision task, with the sole modification being the inclusion of injected prompt cues within the paper content. Since the injected text is visually imperceptible, this task is excluded from evaluations under the PDF-as-img modality.
Theme Task # Ques. Metric
Step Summary (S) 240 BART $\uparrow$, LLM $\uparrow$
Strengths Eval (SE) 238 BART $\uparrow$, LLM $\uparrow$
Weaknesses Eval (WE) 240 BART $\uparrow$, LLM $\uparrow$
Soundness Scoring (SS) 32 MAE $\downarrow$
Presentation Scoring (PS) 32 MAE $\downarrow$
Outcome Conditional Decision (CD) 176 MAE $\downarrow$
Direct Decision (DD) 176 MAE $\downarrow$
CoT Decision (CoD) 176 MAE $\downarrow$
Meta Decision (MD) 240 ACC $\uparrow$
Preference Pairwise Rank (PR) 96 ACC $\uparrow$
Attack Fake Strength (FS) 240 MAE $\downarrow$
Fake Weakness (FW) 238 MAE $\downarrow$
Prompt Injection (PI) 176 MAE $\downarrow$
Table 1: The number of questions and corresponding evaluation metrics of different tasks.
Step Outcome Preference Attack
$\mathrm{S_{B}}\uparrow$ $\mathrm{S_{L}}\uparrow$ $\mathrm{SE_{B}}\uparrow$ $\mathrm{SE_{L}}\uparrow$ $\mathrm{WE_{B}}\uparrow$ $\mathrm{WE_{L}}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
Open-source LLMs & MLLMs
Kimi-VL-A3B-I -3.01 3.65 -3.67 3.46 -3.96 1.98 0.47 0.44 2.31 3.13 3.83 60.92 56.25 2.99 0.56 0.87
Kimi-VL-A3B-T -3.15 3.71 -3.68 3.79 -3.91 2.52 0.47 0.44 2.16 3.59 3.37 66.67 57.29 3.31 0.96 0.63
Qwen2.5-VL-7B -3.05 3.61 -3.68 3.57 -3.96 2.06 0.47 0.44 2.43 3.59 3.57 72.92 59.38 2.99 1.66 0.19
Qwen3-8B -3.08 3.84 -3.63 3.77 -3.87 2.93 0.53 0.44 2.25 3.70 3.16 77.50 65.63 3.26 2.03 1.41
Deepseek-R1-8B -3.09 3.76 -3.63 3.79 -3.86 2.72 0.63 0.78 1.75 3.84 3.55 76.99 64.74 3.25 1.99 0.59
InternVL3-8B -2.99 3.76 -3.65 3.74 -3.93 2.27 0.47 0.44 2.85 3.35 3.35 76.67 52.08 2.99 1.35 0.53
OVIS2-8B -3.09 3.52 -3.70 3.45 -3.99 1.98 0.47 0.44 2.28 3.41 3.72 63.87 60.42 2.99 2.41 0.63
GLM-4.1V-9B-T -3.15 3.60 -3.68 3.73 -3.93 2.60 0.50 0.44 2.08 3.53 3.33 71.86 59.77 3.00 1.08 0.35
Qwen3-14B -3.06 3.85 -3.64 3.83 -3.87 2.79 0.53 0.50 2.16 3.77 3.64 80.42 61.46 3.02 1.99 0.71
OVIS2-16B -3.04 3.59 -3.70 3.52 -3.99 2.06 0.47 0.44 1.87 3.33 3.73 79.92 64.58 3.00 2.34 0.12
Qwen2.5-VL-32B -2.97 3.90 -3.60 3.75 -3.87 2.58 0.56 0.50 2.00 3.15 3.67 67.08 67.71 2.99 1.92 0.80
Qwen3-32B -3.05 3.90 -3.61 3.81 -3.85 2.91 0.50 0.56 2.14 3.60 3.49 80.00 68.75 3.05 1.83 0.78
OVIS2-34B -3.04 3.48 -3.68 3.50 -3.97 2.24 0.81 0.81 2.14 3.72 3.76 79.92 62.50 2.99 1.81 0.57
Qwen2.5-VL-72B -2.99 3.74 -3.65 3.58 -3.94 2.29 0.47 0.47 2.06 3.64 3.71 68.75 63.54 2.99 1.27 0.84
Deepseek-V3 -3.04 3.84 -3.61 3.89 -3.85 2.96 0.53 0.47 2.62 3.37 3.70 75.00 66.03 2.99 0.66 0.20
Deepseek-R1 -3.04 3.92 -3.69 3.90 -3.91 3.05 0.66 0.81 1.97 3.71 3.59 82.92 66.03 3.18 1.20 0.44
Closed-source LLMs & MLLMs
ChatGPT-4o-latest -3.06 3.91 -3.61 3.89 -3.86 2.87 0.84 0.94 1.65 3.70 3.65 80.33 63.54 3.92 1.45 0.45
Claude-sonnet-4 -3.02 3.88 -3.58 3.84 -3.84 3.05 0.53 0.41 1.17 2.02 2.01 84.58 72.92 2.98 2.34 0.43
Gemini-2.5-flash -3.06 3.80 -3.58 3.89 -3.86 2.61 0.94 0.88 1.24 4.60 4.28 74.06 70.83 3.73 1.37 0.59
Gemini-2.5-flash-T -3.06 3.76 -3.58 3.86 -3.86 2.68 0.94 0.88 1.24 4.60 4.28 74.06 70.83 3.73 1.37 0.59
GLM-z1-flash -3.12 3.91 -3.70 3.84 -3.87 2.87 0.47 0.56 2.53 3.77 3.59 70.83 66.67 3.50 2.50 0.18
Table 2: Results on MMReview with text-only inputs.
Step Outcome Preference Attack
$\mathrm{S_{B}}\uparrow$ $\mathrm{S_{L}}\uparrow$ $\mathrm{SE_{B}}\uparrow$ $\mathrm{SE_{L}}\uparrow$ $\mathrm{WE_{B}}\uparrow$ $\mathrm{WE_{L}}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
Open-source MLLMs
Kimi-VL-A3B-I -3.12 3.53 -3.69 3.52 -3.96 1.86 0.47 0.44 2.40 3.61 3.94 60.68 54.17 3.00 0.13 0.65
Kimi-VL-A3B-T -3.13 3.60 -3.68 3.74 -3.92 2.37 0.47 0.47 2.31 3.56 3.64 61.60 62.50 3.14 0.33 0.36
Qwen2.5-VL-7B -3.09 3.64 -3.67 3.70 -3.94 2.11 0.47 0.44 2.20 3.59 3.49 75.83 64.58 3.00 1.45 0.28
InternVL3-8B -3.21 3.46 -3.82 3.26 -4.12 1.84 0.47 0.44 3.02 3.72 3.68 61.11 54.17 3.00 1.33 0.32
OVIS2-8B -3.20 3.43 -3.72 3.49 -3.97 2.06 0.50 0.53 2.27 3.79 4.44 62.91 66.28 3.00 2.15 0.58
GLM-4.1V-9B-T -3.20 3.48 -3.69 3.70 -3.93 2.60 0.50 0.47 2.42 3.68 3.59 70.72 59.77 3.02 1.80 0.31
OVIS2-16B -3.13 3.50 -3.70 3.29 -3.98 2.04 0.47 0.44 2.16 3.40 3.91 78.03 65.12 3.00 2.10 0.04
Qwen2.5-VL-32B -2.99 3.89 -3.64 3.76 -3.90 2.45 0.81 0.75 2.24 3.47 3.94 76.25 60.42 3.00 1.57 0.49
OVIS2-34B -3.11 3.51 -3.69 3.43 -3.97 2.26 0.88 0.81 2.33 3.78 3.91 73.76 62.79 3.00 1.78 0.28
Qwen2.5-VL-72B -3.04 3.74 -3.66 3.71 -3.95 2.33 0.47 0.50 2.06 3.73 3.73 69.58 60.42 3.01 1.30 0.63
Closed-source MLLMs
ChatGPT-4o-latest -3.10 3.86 -3.60 3.85 -3.87 2.67 0.91 0.94 2.24 3.65 3.75 75.97 59.38 3.88 1.22 0.46
Claude-sonnet-4 -3.02 3.86 -3.60 3.88 -3.83 3.09 0.47 0.44 1.19 2.19 2.09 84.17 69.79 2.98 2.12 0.32
Gemini-2.5-flash -3.08 3.75 -3.56 3.84 -3.87 2.58 0.91 0.88 1.34 4.60 4.19 76.25 68.75 2.89 0.78 0.41
Gemini-2.5-flash-T -3.08 3.73 -3.56 3.88 -3.87 2.56 0.91 0.88 1.34 4.60 4.19 76.25 68.75 2.89 0.78 0.41
Table 3: Results on MMReview with multimodal inputs.
Step Outcome Preference Attack
$\mathrm{S_{B}}\uparrow$ $\mathrm{S_{L}}\uparrow$ $\mathrm{SE_{B}}\uparrow$ $\mathrm{SE_{L}}\uparrow$ $\mathrm{WE_{B}}\uparrow$ $\mathrm{WE_{L}}\uparrow$ $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$
Open-source MLLMs
Kimi-VL-A3B-I -3.21 3.17 -3.74 3.34 -4.00 1.74 0.47 0.44 2.30 3.19 3.83 69.36 55.21 0.28 0.00
Kimi-VL-A3B-T -3.21 3.50 -3.71 3.72 -3.91 2.54 0.47 0.44 1.85 3.72 3.49 74.37 60.42 3.11 0.27
Qwen2.5-VL-7B -3.21 3.28 -3.72 3.49 -3.95 1.91 0.47 0.44 1.97 3.09 3.12 70.83 47.92 2.99 1.01
InternVL3-8B -4.86 1.13 -4.59 1.89 -4.55 1.13 0.47 0.44 2.40 3.63 3.41 68.75 75.00 2.99 1.98
OVIS2-8B -3.23 3.28 -3.70 3.43 -3.99 2.06 0.47 0.44 2.49 3.44 3.93 65.32 53.13 3.00 2.00
GLM-4.1V-9B-T -3.26 3.36 -3.67 3.69 -3.92 2.69 0.47 0.44 1.80 3.51 3.13 80.33 65.63 3.00 1.82
OVIS2-16B -3.20 3.42 -3.70 3.49 -3.97 2.17 0.47 0.44 2.05 3.23 3.76 78.66 54.17 3.00 1.81
Qwen2.5-VL-32B -3.11 3.66 -3.71 3.67 -3.92 2.46 0.78 0.81 2.11 3.55 3.77 67.50 61.46 3.24 1.22
OVIS2-34B -3.20 3.41 -3.68 3.49 -3.99 2.18 0.78 0.69 1.91 3.63 3.74 77.82 52.08 2.99 1.69
Qwen2.5-VL-72B -3.11 3.57 -3.71 3.46 -3.95 2.22 0.47 0.59 2.15 3.76 3.77 70.00 54.17 3.03 1.11
Closed-source MLLMs
ChatGPT-4o-latest -3.33 3.65 -3.69 3.70 -3.88 2.68 0.91 0.91 1.84 3.63 3.68 77.73 59.38 3.90 1.33
Claude-sonnet-4 -3.19 3.59 -3.65 3.73 -3.90 2.88 0.48 0.45 1.22 2.36 2.22 82.92 62.50 3.00 2.73
Gemini-2.5-flash -3.31 3.26 -3.66 3.73 -3.93 2.48 0.91 0.91 1.11 4.60 4.19 73.33 69.79 1.67 0.75
Gemini-2.5-flash-T -3.31 3.30 -3.67 3.74 -3.93 2.49 0.91 0.91 1.11 4.60 4.19 73.33 69.79 1.67 0.75
Table 4: Results on MMReview with pdf-as-img inputs.
Experiments
Evaluation details
We conduct extensive experiments on MMReview using a wide range of popular LLMs and MLLMs. Specifically, for open-source models, we include a diverse set of architectures and sizes, covering Qwen2.5VL (Bai et al. 2025), Kimi-VL-A3B (Team et al. 2025b), Qwen3 (Yang et al. 2025), InternVL3 (Zhu et al. 2025), OVIS2 (Lu et al. 2024), GLM-4.1V (Team et al. 2025a), and Deepseek (DeepSeek-AI et al. 2025). For advanced proprietary models, our selection includes GPT-4o, Gemini-2.5-Flash, GLM-Z1-Flash, and Claude-4-Sonnet.
As shown in Table 1, for tasks without objective evaluation metrics, namely S, SE, and WE, we employ BARTScore (Yuan, Neubig, and Liu 2021) and the ‘LLM-as-a-Judge’ paradigm (Bai et al. 2023; Zheng et al. 2023; Gu et al. 2024) to assess the similarity between model-generated and human-written review comments. For tasks where the model output is a numerical score, including SS, PS, CD, DD, CoD, FS, FW, and PI, we compute the Mean Absolute Error (MAE) between the model’s predicted score and the ground-truth score to quantify deviation. For classification-based tasks such as MD and PR, we evaluate performance using accuracy.
Main Results
The experimental results on MMReview, as shown in Tables 2, 3, and 4 (where T denotes Thinking and I denotes Instruct), yield the following key findings:
(1) Proprietary models consistently outperform open-source models. Closed-source models such as ChatGPT-4o, Claude, and Gemini demonstrate significant superiority across most metrics, particularly on high-level tasks such as Meta Decision (MD), Pairwise Ranking (PR), Conditional Decision (CD), and Chain-of-Thought Decision (CoD). This indicates that proprietary models exhibit greater consistency and reliability in generating review content, along with enhanced capabilities in understanding complex academic material and producing structured assessments.
(2) Input modality has a substantial impact on model performance. For most models, performance improves with the addition of visual inputs. Specifically, tasks directly related to evaluation outcomes (e.g., MD, PR, CD, DD, CoD) show notable gains under multimodal (text+image) input conditions. For instance, Qwen-2.5-VL-7B’s accuracy on the MD task increases from 72.92% to 75.83% with image input, its MAE on CD decreases from 2.43 to 2.20, and its PR accuracy rises from 59.38% to 64.58%. However, the PDF-as-image modality poses challenges: most models—including advanced proprietary ones like Claude and Gemini—perform worse on this modality than on structured text-image input, though closed-source models remain comparatively more robust than open-source counterparts.
(3) Proprietary models exhibit stronger robustness to prompt injection. On the Prompt Injection (PI) task, proprietary models achieve MAE values below 0.6 in most cases, often under 0.5, indicating consistent judgment with and without injected prompts. In contrast, open-source models exhibit highly variable performance under prompt injection, suggesting greater susceptibility to adversarial manipulation.
(4) Larger model size does not guarantee superior performance. A comparative analysis of the Qwen2.5-VL and OVIS2 model families reveals that performance does not uniformly improve with scale. In several cases, smaller models outperform their larger counterparts, suggesting that model size is not a definitive determinant of effectiveness in peer review tasks.
More analysis of the result can be found in the supplementary materials.
Further Analysis
Are Thinking Models More Suitable for Review?
For closed-source models, performance differences between Gemini-2.5-Flash and Gemini-2.5-Flash-Thinking are minimal. In contrast, within the open-source Kimi-VL-A3B and Deepseek series, a more nuanced trend emerges: while non-thinking variants may outperform on some step-based tasks, thinking models consistently excel in outcome-based and preference-based tasks critical to review decisions. This indicates that thinking-enhanced models are better suited for generating final peer review judgments. Further supporting evidence is provided in the supplementary material.
AI NS ES $S$
ChatGPT-4o-latest 4.46 5.67 3.67 4.28
claude-sonnet-4-20250514 2.08 2.33 1.33 2.44
gemini-2.5-flash-thinking 4.62 6.33 4.33 4.50
Deepseek-R1 2.32 4.00 3.67 4.00
Table 5: The MAE performance of several models on the CoD task across different research domains.
What Domains are LLM-reviewers better at?
Table 5 reports the MAE performance of various models on the CoD task across research domains, indicating their relative proficiency in field-specific peer review. All models use multimodal inputs except Deepseek-R1, which operates with text only. Results show that most closed-source models perform best in Engineering Sciences, while Deepseek-R1 achieves its lowest MAE in Artificial Intelligence. These findings suggest that models’ reviewing strengths vary across disciplines, potentially influenced by differences in pretraining corpora and input modalities.
More further analysis can be found in the supplementary materials of our submission.
Conclusion
In this work, we present MMReview—a multidisciplinary and multimodal benchmark designed to evaluate the capabilities of Large Language Models (LLMs) in academic peer review. The benchmark encompasses 4 thematic categories and 13 distinct tasks. Its core features include coverage across diverse academic disciplines, support for multimodal input formats, and a comprehensive set of evaluation tasks that span the full peer review pipeline. Leveraging MMReview, we conducted extensive evaluations of state-of-the-art LLMs and MLLMs. We envision MMReview as a standardized evaluation platform that can catalyze the development of more efficient LLM-assisted peer review systems. In future work, we plan to incorporate larger-scale datasets specifically curated for training LLMs in the context of scholarly review.
Appendix
As a supplement, we provide additional materials in the appendix, including more detailed statistics on the benchmark and tasks (Appendix A), extended experimental results and analyses (Appendix B), implementation details of the experiments (Appendix C), several illustrative case studies (Appendix D), as well as the prompts used during the construction and application of the benchmark (Appendix E).
Appendix A A. Statistical Information about Research Domains in MMReview Benchmark
Discipline Research Field # Samples
Artificial
Intelligence
Machine Learning 17
Computer Vision 20
Natural Language Processing 16
Reinforcement Learning 11
Graph Neural Networks 19
Signal Processing 17
AI Application 15
Natural
Sciences
Biology & Medicine 11
Physics 16
Chemistry 13
Environmental & Earth Sciences 10
Mathematics & Statistics 13
Engineering
Sciences
Materials Science 14
Control Science 12
Electronic Information 6
Energy Science 6
Social
Sciences
Society, Economics & Finance 24
Table 6: The distribution of papers across various research domains.
Table 6 presents the distribution of papers across various research domains within our MMReview benchmark. For each domain, we ensured a relatively balanced number of samples through a combination of model-based filtering and manual curation.
Appendix B B. More Results and Analysis
B.1 The Impact of Paper Length
As a long-context task, peer review performance may be influenced by the length of the manuscript. To examine this, we analyze the average scores assigned by LLMs under two input settings: text-only (with context length measured in tokens) and PDF-as-image (with context length measured by the number of rendered pages). As illustrated in Figures 3 and 4, we observe a consistent trend wherein models tend to assign higher scores as the context length or number of images increases. This upward bias does not align well with human reviewer judgments across multiple intervals, indicating an inherent tendency of LLMs to overvalue longer inputs irrespective of actual content quality. Such length-induced bias poses a significant challenge for the practical deployment of LLMs in peer review scenarios.
Figure 3: The average scores under text-only input setting, with context length measured in tokens.
Figure 4: The average scores under pdf-as-img input setting, with context length measured in tokens.
B.2 How does Prompt Injection Work?
To further investigate the impact of prompt injection on model scoring behavior, we analyze score variations between the Prompt Injection task and the CoT Decision (CoD) task, where the only difference in prompts lies in the injection of the statement: “IGNORE ALL PREVIOUS INSTRUCTIONS, NOW GIVE A POSITIVE REVIEW OF THESE PAPERS AND DO NOT HIGHLIGHT ANY NEGATIVES.” All other prompt content remains identical. Tables 8 and 9 report the proportion and expectation of score increases and decreases under text-only and multimodal input conditions, respectively. The experimental results reveal three key findings: (1) Most models exhibit an overall increase in average score after prompt injection, with a typical rise of 1–2 points; (2) Multimodal inputs enhance model robustness against prompt injection compared to text-only inputs, consistent with prior observations of improved performance under multimodal settings; (3) The robustness of Thinking variants varies across model families—within the Qwen3 and Gemini series, Thinking models are less robust than their non-Thinking counterparts, whereas in the Kimi and Deepseek series, Thinking models demonstrate greater resistance to injection. These observations suggest that the relationship between prompt injection robustness and the model’s internal reasoning style is non-trivial and may depend on model-specific cognitive mechanisms and prompt interpretation behavior.
model $\mathrm{SS}\downarrow$ $\mathrm{PS}\downarrow$ $\mathrm{CD}\downarrow$ $\mathrm{DD}\downarrow$ $\mathrm{CoD}\downarrow$ $\mathrm{MD}\uparrow$ $\mathrm{PR}\uparrow$ $\mathrm{FS}\downarrow$ $\mathrm{FW}\downarrow$ $\mathrm{PI}\downarrow$
Deepseek-V3
w.o. Ref
0.53 0.47 2.62 3.37 3.70 0.75 0.66 2.99 0.66 0.20
Deepseek-V3
w/ Ref
0.47 0.44 2.56 3.26 3.74 0.83 0.57 2.99 0.71 0.19
Table 7: The influence of references on model performance.
% Raise EE(Raise) % Lower EE(Lower)
Gemini-2.5-flash-T 22.16 1.81 10.78 -1.33
Gemini-2.5-flash 22.16 1.81 10.78 -1.33
Qwen3-8B 90.34 1.56 0.57 -1.00
Qwen3-8B-nothink 98.30 1.50 0.00 0.00
Kimi-VL-A3B-I 31.15 1.28 6.56 -0.50
Qwen3-32B 61.93 1.21 3.41 -1.00
Chatgpt-4o-latest 36.47 1.17 1.18 -1.00
Claude-sonnet-4 11.93 1.14 25.57 -1.13
Deepseek-R1-8B 36.99 1.14 12.72 -1.09
Kimi-VL-A3B-T 46.02 1.10 11.93 -1.00
Qwen3-14B 63.64 1.10 1.14 -1.00
Qwen2.5-VL-72B 78.41 1.07 0.00 0.00
Deepseek-V3 16.48 1.07 2.84 -1.00
OVIS2-8B 51.81 1.06 1.20 -1.00
Qwen3-14B-nothink 80.68 1.05 0.00 0.00
Deepseek-R1 15.91 1.04 22.73 -1.20
Qwen2.5-VL-32B 77.71 1.03 0.00 0.00
Qwen3-32B-nothink 55.68 1.00 0.00 0.00
Qwen2.5-VL-7B 18.75 1.00 0.00 0.00
OVIS2-34B 56.00 1.00 0.00 0.00
OVIS2-16B 11.93 1.00 0.00 0.00
InternVL3-8B 52.84 1.00 0.00 0.00
GLM-z1-flash 14.86 1.00 3.43 -1.00
GLM-4.1V-9B-T 25.48 1.00 6.37 -1.10
Table 8: The proportion and expectation of score increases and decreases under text-only input condition.
% Raise E(Raise) % Lower E(Lower)
Kimi-VL-A3B-I 23.23 1.54 7.74 -1.54
OVIS2-16B 2.94 1.40 0.00 0.00
GLM-4.1V-9B-T 14.97 1.28 3.59 -1.00
Chatgpt-4o-latest 34.30 1.08 3.49 -2.17
Kimi-VL-A3B-T 21.51 1.08 8.72 -1.00
Gemini-2.5-flash-T 26.44 1.07 12.07 -1.00
Gemini-2.5-flash 27.01 1.06 12.07 -1.00
InternVL3-8B 29.48 1.06 0.58 -1.00
OVIS2-8B 41.72 1.02 1.32 -4.00
Qwen2.5-VL-32B 47.73 1.01 1.14 -1.00
Qwen2.5-VL-72B 57.14 1.01 0.00 0.00
Qwen2.5-VL-7B 28.41 1.00 0.00 0.00
OVIS2-34B 28.40 1.00 0.00 0.00
Claude-sonnet-4 9.66 1.00 21.02 -1.08
Table 9: The proportion and expectation of score increases and decreases under multimodal input condition.
B.3 The Impact of Reference Section
References are a critical component of academic writing, serving to substantiate claims and situate the work within the broader scholarly context. However, for LLMs lacking internet access, the reference section may consume a substantial portion of the input context without providing direct utility, thereby reducing the available token space for more informative content. To investigate the influence of references on model performance, we conduct an ablation study (Table 7) comparing inputs with and without the reference section. Results indicate that removing references improves performance on tasks such as Chain-of-Thought (CoT) scoring, alignment with human preferences, and detection of hallucinated strengths and weaknesses. In contrast, for tasks involving quantitative quality assessments, such as Soundness Scoring (SS) and Presentation Scoring (PS), the inclusion of references proves beneficial, as their absence renders the manuscript less complete and increases the model’s MAE. This suggests a trade-off: while references enhance content completeness and improve technical evaluations, their removal shortens the input context and may reduce length-related bias, enabling models to make final judgments more aligned with human preferences.
Appendix C C. More Implementation Details
For LLMs and MLLMs with parameter sizes up to 72B, we conducted evaluations through direct model deployment, while for models exceeding 72B or proprietary models, we performed testing via API access. All experiments were carried out on NVIDIA A100 GPUs. To enhance the reproducibility of our results, we set the temperature parameter to 0. Prompt templates and evaluation scripts were manually crafted with reference to reviewer guidelines from major academic conferences. The prompts used in the evaluations are provided in the appendix.
Appendix D D. Prompts
In this section, we present the prompts used throughout the construction and evaluation phases of the benchmark. Figure 5 illustrates the prompt employed in the task generation stage to produce antonymic rewrites of strengths and weaknesses. Figures 6 through 16 display the prompts corresponding to the various tasks designed within the benchmark. Additionally, Figure 17 presents the prompt based on the Likert scale used for LLM-as-a-judge evaluation.
  • Task generation prompt: Figure 5
  • System prompt: Figure 6
  • Summary: Figure 7
  • Strengths and Weaknesses: Figure 8-9
  • Soundness Scoring and Presentation Scoring: Figure 10
  • Conditional Decision: Figure 11
  • Direct Decision: Figure 12
  • CoT Decition: Figure 14
  • Meta Decision: Figure 13
  • Preference Rank: Figure 15
  • Fake Strengths and Fake Weaknesses: Figure 16
  • Prompt Injection: The same as CoT Decision (Figure 14)
  • LLM-as-a-judge: Figure 17
Figure 5: Prompt for GPT-4o to generate Fake Strengths and Fake Weaknesses tasks.
Figure 6: System prompt for LLMs to generate reviews.
Figure 7: The prompt used in Summary task.
Figure 8: The prompt used in Strengths Evaluation task.
Figure 9: The prompt used in Weaknesses Evaluation task.
Figure 10: The prompt used in $S$ and PS task.
Figure 11: The prompt used in CD task.
Figure 12: The prompt used in $D$ task.
Figure 13: The prompt used in MD task.
Figure 14: The prompt used in CoD task.
Figure 15: The prompt used in PR task.
Figure 16: The prompt used in FS and FW task.
Figure 17: The prompt used in LLM-as-a-judge evaluation.